OpenStudy (anonymous):

Find two real numbers whose sum is 4 and whose product is a maximum.

OpenStudy (anonymous):

so maximize xz where x + z = 4....{ x = z = 2 }.. f(x) = x [ 4 -x] is a parabola...max easy to find

OpenStudy (anonymous):

-------------- Sum: x + y = 4 Product = xy ------------------- y = 4-x Substitute that into "Product" to get: P = x(4-x) P = 4x-x^2 That is a quadratic with a = -1, b = 4 ---- Comment: That "Product" does not have a minimum; it does have a maximum. Max occurs at x = -b/2a = -4/(2*-1) = 2 Maximum product is P = 2*(4-2)= 2*2 = 4

OpenStudy (agent0smith):

You can solve this easily without any algebra... both numbers must be positive for them to add up to 4 *and* have a maximum product... (if one is negative and the other positive, the product is negative) Numbers that add to 4: 0 and 4 1 and 3 2 and 2 and those are your only options. Find the two that have the greatest product.